# No Free Prune: Information-Theoretic Barriers to Pruning at Initialization

### v1

@@ B.1.3 Isoperimetry, Subgaussianity, and Crude Bounds
where we highlight the explicit dependence on $C_{0}$. Thus Lemma 2 yields
@@
$a_{1}\sqrt{C_{0}\left(\frac{I(\{X_{t}\}_{t\in T};W)}{\delta}+\log\frac{2}{delta}\right)}\geq\frac{\epsilon}{8}$
@@
B.1.4 Controlling Mutual Information
We next show Lemma 3, which gives a upper bound for $I(\{X_{t}\}_{t\in T};W)$.
Proof of Lemma 3.
@@ B.1.4 Controlling Mutual Information
Combining Lemma 6 with equation (7), followed by applying Lemma 3 then yields the desired Theorem 3.
B.2 The continuous setting
@@
Define $\mathcal{W}_{\mathbf{m},L,\epsilon}$ to be an $\frac{\epsilon}{8J}$-net of $\mathcal{W}_{\mathbf{m},L}$. Note that since $\mathrm{diam}(\mathcal{W})\leq W$ and $\dim\mathcal{W}_{\mathbf{m},L}\leq\|\mathbf{m}\|_{1}$, that $N_{\mathbf{m},L,\epsilon}=|\mathcal{W}_{\mathbf{m},L,\epsilon}|\leq(1+60WJ\epsilon^{-1})^{\|\mathbf{m}\|_{1}}$ (Corollary 4.2.13 of Vershynin (2018)). We then apply Lemma 6 to
@@
We are now in a position to apply Lemma 6, obtaining
@@
Then with probability $1-\delta$ over the data, $f^{W}$ must have Lipschitz constant $\operatorname{Lip}(f)\geq L$. Finally, by Lemma 3, we have
@@
with probability $1-\delta$, concluding the proof of Theorem 4.
Appendix C Experimental Details
C.1 Experimental Details
C.1.1 Figures 0(a) and 0(b)
@@
For FashionMNIST, we added $\sigma^{2}=3$ Gaussian noise to each input image, preserving its original label, allowing us to see if our predictions hold in general settings beyond label corruption and finding, as expected, indeed they do. We train till convergence in loss to within $0.01$ (or until accuracy doesn’t change for three consecutive epochs), with $\eta=1e-3$ on Adam. We use a batch size of $64$ with a two-hidden layer ReLU architecture with a hidden width of $200$. We plotted the mean and standard error (standard deviation of the mean) over $k=10$ different seeds on this dataset. This amount of repetitions with standard datasets like MNIST is common in the literature; Frankle et al. (2020b) use 5 repetitions on similar datasets like CIFAR10, and less for larger datasets like ImageNet.
We used the pruning code of Tanaka et al. (2020) to guarantee consistency in implementation of pruning methods with the rest of the literature. This includes their implementation of all pruning algorithms, though we made modifications to, for instance, add magnitude pruning after training, which they do not include. For the Gaussian case, since the infrastructure for it was not in Tanaka et al. (2020), we wrote our own implementations of the pruning algorithms based on the papers where they were introduced, and checked our implementations’ outputs against those of Tanaka et al. (2020) where possible, finding agreement. There is further discussion around implementation details of pruning algorithms in Frankle et al. (2020b), where they note that their plots mostly agree with those of Tanaka et al. (2020), with some small differences.
C.1.2 Figures 1(a), 1(b), 1(c)
@@
In the case of figures 1(a), we track this quantity over the course of IMP on a student-teacher task with noise scale $\sigma^{2}=1$. Changing the noise scale simply rescales the y-axis without affecting trend we see. We use a 5-layer ReLU MLP student with hidden width $100$ and a 3-layer teacher with hidden width $50$. This mismatch reflects the fact that our architecture rarely reflects the data-generating process perfectly. We train with Adam with $\eta=1e-3$, pruning the lowest $20\%$ of weights out at every iteration of IMP, as is standard Frankle & Carbin (2018); Frankle et al. (2019). We use $n=1000$ data points of dimension $d=50$, where both the student and teacher produce scalar outputs. We average and plot the mean and standard error over different random seeds generating $k=25$ networks (student and teacher) and datasets. The standard error is so small in the plot it is barely visibly. Then for Figure 1(b), 1(c), we use the same setup as above except do not prune, and sweep over learning rates and gradient noise instead, tracking the same quantities. These are also averaged over $k=25$ random seeds.
C.1.3 Tail end of Figure 1(a)
We see that Figure 1(a) gives the expected trend of capacity increasing and decreasing as we train, then prune iteratively during the IMP algorithm to find lottery tickets. Of course, one cannot keep pruning every $100$ epochs forever; eventually one reaches an empty network, so an important sanity check is that such ability to correlate decays at late-time as the network is pruned completely and all weights vanish. We can see below that effective parameter count goes becomes low at extreme sparsities when the network has almost no actual parameters remaining, as one might expect. Again, this is averaged over $k=25$ networks. The critical sparsity level at which this correlation ability begins to decay is the lottery ticket sparsity. For us this occurs after $\approx 9$ pruning iterations where we drop $20\%$ of the weights each time, at which point only $\sim 10\%$ of the weights remain.
Figure 3: Correlation with data noise over training epochs, IMP. Expanded version of 1(a) pruned to almost complete sparsity.
C.2 Toy model: Computing MI exactly
@@
Appendix D Generalization Bounds
D.1 Information Theoretic Generalization Bounds
By applying Lemma 1, one can find analogs of effective parameter count in classical generalization bounds. Namely, a low mutual information choice of function class with good generalization will also have good generalization, even if the number of such function classes is very large. To begin, define the generalization error for a function $h$ on a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ of i.i.d input-output pairs in $\mathcal{X}\times\mathcal{Y}$ with respect to a loss $\ell$ as
@@ D.1 Information Theoretic Generalization Bounds
Theorem 5 (VC Dimension Generalization Bound).
@@
Recall that $n$-sample VC dimension generalization bounds for a function class of VC dimension $d$ scale roughly as $\tilde{O}\left(\sqrt{\frac{d}{n}}\right)$. The bound above instead scales as $\tilde{O}\left(\sqrt{\frac{d+\delta^{-1}I}{2m}}\right)$. Although we now have worse dependence on $\delta$ (from Lemma 1), if one views $\delta$ as a small constant, we recover the same asymptotics with $d+I$ in place of $d$. This is an analog of our main result if one interprets $d+I$ as an “effective VC dimension”.
@@
$\mathrm{Rad}_{\mathcal{D}}(\mathcal{H})=\frac{1}{m}\mathbb{E}_{\bm{\sigma}\sim\{\pm 1\}^{m}}\left[\sup_{h\in\mathcal{H}}\sum_{i=1}^{m}\sigma_{i}h(x_{i})\right]$ (11)
Theorem 6.
@@
The above bound again illustrates that incorporating mutual information into the choice of function class yields generalization bounds which decay gracefully with the value of $I$. In the case that each function class has at most $N$ functions, applying Massart’s lemma Shalev-Shwartz & Ben-David (2014) yields
@@
D.2 Proof of Theorem 5 (VC Dimension)
@@
Here the first inequality is Theorem 6.11 from Shalev-Shwartz & Ben-David (2014) while the second is Sauer’s lemma. Our strategy will be to argue that the bound above holds with high probability for fixed $\mathcal{H}$ by the bounded differences inequality and conclude using Lemma 1.
Proof of Theorem 5.
@@
i.e. $X_{t}$ is $(1/2m)$-subgaussian. Applying Lemma 1 now implies that with probability $1-\delta$, one has
@@
D.3 Proof of Theorem 6 (Rademacher Complexity)
@@
This is Lemma 26.2 of Shalev-Shwartz & Ben-David (2014). Our strategy is equivalent to the above.
@@
We once more define $Y_{t}=\sup_{h\in\mathcal{H}_{t}}\mathrm{gen}_{\mathcal{D}}(h)$ and $X_{t}=Y_{t}-\mathbb{E}[Y_{t}]$. Note that $X_{t}$ changes by at most $2b/m$ when a given datapoint is changed, and thus McDiarmid’s bounded differences inequality again implies that $X_{t}$ is $(2b^{2}/m)$-subgaussian. Then applying Lemma 1 yields that with probability $1-\delta$,
@@
Applying Lemma 26.2 Shalev-Shwartz & Ben-David (2014) above, we obtain the following almost sure inequality:
@@
We again note that $\mathrm{Rad}_{\mathcal{D}}(\ell\circ\mathcal{H}_{W})$ changes by at most $2b/m$ when changing an individual datapoint. Hence, once centered, it is again $(2b^{2}/m)$-subgaussian. Applying Lemma 1 again produces that with probability $1-\delta$,
@@
Noting that $\ell$ is Lipschitz in its first argument, we apply Talagrand’s contraction lemma (see e.g. Ledoux & Talagrand (2013); Shalev-Shwartz & Ben-David (2014)) to conclude the proof of Theorem 6. ∎

### v2

@@ A.1.3 Isoperimetry, Subgaussianity, and Crude Bounds
where we highlight the explicit dependence on $C_{0}$. Thus Lemma 5.4 yields
@@
$a_{1}\sqrt{C_{0}\left(\frac{I(\{X_{t}\}_{t\in T};W)}{\delta}+\log\frac{2}{\delta}\right)}\geq\frac{\epsilon}{8}$
@@
A.1.4 Controlling Mutual Information
We next show Lemma 5.5, which gives a upper bound for $I(\{X_{t}\}_{t\in T};W)$.
Proof of Lemma 5.5.
@@ A.1.4 Controlling Mutual Information
Combining Lemma A.3 with equation (9), followed by applying Lemma 5.5 then yields the desired Theorem 5.3.
A.2 The continuous setting
@@
Define $\mathcal{W}_{\mathbf{m},L,\epsilon}$ to be an $\frac{\epsilon}{8J}$-net of $\mathcal{W}_{\mathbf{m},L}$. Note that since $\mathrm{diam}(\mathcal{W})\leq W$ and $\dim\mathcal{W}_{\mathbf{m},L}\leq\|\mathbf{m}\|_{1}$, that $N_{\mathbf{m},L,\epsilon}=|\mathcal{W}_{\mathbf{m},L,\epsilon}|\leq(1+60WJ\epsilon^{-1})^{\|\mathbf{m}\|_{1}}$ (Corollary 4.2.13 of (Vershynin, 2018)). We then apply Lemma A.3 to
@@
We are now in a position to apply Lemma A.3, obtaining
@@
Then with probability $1-\delta$ over the data, $f^{W}$ must have Lipschitz constant $\operatorname{Lip}(f)\geq L$. Finally, by Lemma 5.5, we have
@@
with probability $1-\delta$, concluding the proof of Theorem 5.6.
Appendix B Experimental Details
B.1 Experimental Details
B.1.1 Figures 1(a) and 1(b)
@@
For FashionMNIST, we added $\sigma^{2}=3$ Gaussian noise to each input image, preserving its original label, allowing us to see if our predictions hold in general settings beyond label corruption and finding, as expected, indeed they do. We train till convergence in loss to within $0.01$ (or until accuracy doesn’t change for three consecutive epochs), with $\eta=1e-3$ on Adam. We use a batch size of $64$ with a two-hidden layer ReLU architecture with a hidden width of $200$. We plotted the mean and standard error (standard deviation of the mean) over $k=10$ different seeds on this dataset. This amount of repetitions with standard datasets like MNIST is common in the literature; (Frankle et al., 2020b) use 5 repetitions on similar datasets like CIFAR10, and less for larger datasets like ImageNet.
We used the pruning code of (Tanaka et al., 2020) to guarantee consistency in implementation of pruning methods with the rest of the literature. This includes their implementation of all pruning algorithms, though we made modifications to, for instance, add magnitude pruning after training, which they do not include. For the Gaussian case, since the infrastructure for it was not in (Tanaka et al., 2020), we wrote our own implementations of the pruning algorithms based on the papers where they were introduced, and checked our implementations’ outputs against those of (Tanaka et al., 2020) where possible, finding agreement. There is further discussion around implementation details of pruning algorithms in (Frankle et al., 2020b), where they note that their plots mostly agree with those of (Tanaka et al., 2020), with some small differences.
B.1.2 Figures 1, 1, 1
@@
In the case of figures 1, we track this quantity over the course of IMP on a student-teacher task with noise scale $\sigma^{2}=1$. Changing the noise scale simply rescales the y-axis without affecting trend we see. We use a 5-layer ReLU MLP student with hidden width $100$ and a 3-layer teacher with hidden width $50$. This mismatch reflects the fact that our architecture rarely reflects the data-generating process perfectly. We train with Adam with $\eta=1e-3$, pruning the lowest $20\%$ of weights out at every iteration of IMP, as is standard (Frankle & Carbin, 2018; Frankle et al., 2019). We use $n=1000$ data points of dimension $d=50$, where both the student and teacher produce scalar outputs. We average and plot the mean and standard error over different random seeds generating $k=25$ networks (student and teacher) and datasets. The standard error is so small in the plot it is barely visibly. Then for Figure 1, 1, we use the same setup as above except do not prune, and sweep over learning rates and gradient noise instead, tracking the same quantities. These are also averaged over $k=25$ random seeds.
B.1.3 Tail end of Figure 1
We see that Figure 1 gives the expected trend of capacity increasing and decreasing as we train, then prune iteratively during the IMP algorithm to find lottery tickets. Of course, one cannot keep pruning every $100$ epochs forever; eventually one reaches an empty network, so an important sanity check is that such ability to correlate decays at late-time as the network is pruned completely and all weights vanish. We can see below that effective parameter count goes becomes low at extreme sparsities when the network has almost no actual parameters remaining, as one might expect. Again, this is averaged over $k=25$ networks. The critical sparsity level at which this correlation ability begins to decay is the lottery ticket sparsity. For us this occurs after $\approx 9$ pruning iterations where we drop $20\%$ of the weights each time, at which point only $\sim 10\%$ of the weights remain.
Figure 2: Correlation with data noise over training epochs, IMP. Expanded version of 1 pruned to almost complete sparsity.
B.2 Toy model: Computing MI exactly
@@
Appendix C Generalization Bounds
C.1 Understanding the Information-Parameter Tradeoff
Here, we elaborate on the general principle at the heart of our main result: that for the sake of many theoretical generalization guarantees, parameter count of the model and the model’s mutual information with the data are interchangeable. We illustrate this in a toy model.
C.1.1 A saturating example
First, we present a toy example which saturates our bound Theorem 5.6.
We take $n=d^{C}$ for some constant $C$. For the sake of illustration, consider a dataset of $n$ covariates $x_{1},\dots,x_{n}\stackrel{{\scriptstyle IID}}{{\sim}}\mathsf{N}(0,I_{d}/d)$ in dimension $d$, with uniformly random labels $y_{1},\dots,y_{n}\in\{\pm 1\}$. We now consider the following masked function class. First, we sample a single, large matrix with i.i.d. standard Gaussian entries $\mathbf{W}\in\mathbb{R}^{p\times d}$. Next, we construct the following function class:
=$\mathcal{F}=\{f(x)=(\mathbf{a}\cdot\mathbf{m})^{\top}\mathrm{ReLU}(\mathbf{W}x+\mathbf{b})\mid\mathbf{m}\in\{0,1\}^{p}\}.$ (12)
We will take $\mathbf{b}\in\mathbb{R}^{p}$ to have all coordinates $-0.8$. First let us point out that $\mathcal{F}$ is a one-hidden layer neural network with ReLU activations, and width $p$. The only learnable parameters of this model are in the choice of mask $\mathbf{m}$ the weights are randomly drawn at the start and fixed. Hence this model has $p$ parameters. The mask entries determines whether a certain hidden layer node is left nonzero, or is forcibly masked away to zero.
We now consider $p=\exp(O(nd))$. The idea is now that for this toy model, each $x_{i}$, of which there are $n$, can be memorized by a neuron. Since there are exponentially many hidden nodes, each observing the one dimensional projection $w_{j}^{\top}x$, there exists, with high probability, one that aligns very closely with $x_{i}$ in the sense that $\|w_{j}-x\|\leq\|x\|/100$. Call this one $w_{\text{mem}(i)}$, for the memorizer of $i$. Since $n=d^{C}$, a standard union bound shows that with high probability, for all $i,j$:
$\|x_{i}\|$ $\in[0.99,1.01],$
$|\langle x_{i},x_{j}\rangle|$ $\leq d^{-1/3}.$
The triangle inequality then implies (assuming the preceding estimates hold) that $\langle w_{\text{mem}(i)},x_{i}\rangle\in[0.9,1.1]$ and
$|\langle w_{\text{mem}(i)},x_{j}\rangle|\leq 1/10,\quad\forall j\neq i.$
Thus, we can set the value of $\mathbf{b}$ to be $-0.8$ in all entries, so that $w^{\top}_{\text{mem}(i)}x_{i}+b_{\text{mem}(i)}$ is positive only for the $i$-th covariate, and negative for all the others. After applying ReLU, it is only nonzero for the $i$-th sample – hence, the $(\text{mem}(i))$-th hidden node indeed memorizes the $i$-th sample.
Now consider the pruning algorithm which looks at the data and sets the mask to only be nonzero on the nodes $\{\text{mem}(i)\mid i\in[n]\}$. This function is $O(1)$-Lipschitz on $\{x\in\mathbb{R}^{d}~{}:~{}\|x\|\leq 1.1\}$. This is because such any such $x$ satisfies $\langle x,w_{\text{mem}(i)}\rangle\geq 0.8$ for at most $1$ value of $i$. This follows easily from the fact that different $w_{\text{mem}(i)}\approx x_{i}$ are approximately orthogonal, and $1.280.8^{2}+0.8^{2}=$ 1.28 is significantly larger than $1.1^{2}$.
Moreover this function has only $s=n\ll nd$ nonzero parameters. However it clearly has $\Theta(nd)$ mutual information with the data ($\Theta(d)$ bits per sample); hence it has $\Theta(nd)$ effective parameters. Thus, this presents a scenario saturating our bound (Theorem 5.6). In some sense, these are “lottery tickets” - if we trained this network, we expect a weight $w_{k}$ with high alignment to a given input to have high learned coefficient $a_{k}$; our mask simply selectively chooses a single node with high alignment for each sample.
More generally, Theorem 5.6 implies that if one wants $O(nd)$ effective parameters with $O(1)$ Lipschitz constant, while maintaining $s\ll O(nd)$ nonzero parameters, then $\Omega(nd)$ mutual information with the data is required. This means that, including the mask, one actually needs $\exp(\Omega(nd/s))$ total parameters in the model (c.f. (Bubeck et al., 2021), Theorem 4).
C.2 Information Theoretic Generalization Bounds
By applying Lemma 5.2, one can find analogs of effective parameter count in classical generalization bounds. Namely, a low mutual information choice of function class with good generalization will also have good generalization, even if the number of such function classes is very large. To begin, define the generalization error for a function $h$ on a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ of i.i.d input-output pairs in $\mathcal{X}\times\mathcal{Y}$ with respect to a loss $\ell$ as
@@ C.2 Information Theoretic Generalization Bounds
Theorem C.1 (VC Dimension Generalization Bound).
@@
Recall that $n$-sample VC dimension generalization bounds for a function class of VC dimension $d$ scale roughly as $\tilde{O}\left(\sqrt{\frac{d}{n}}\right)$. The bound above instead scales as $\tilde{O}\left(\sqrt{\frac{d+\delta^{-1}I}{2m}}\right)$. Although we now have worse dependence on $\delta$ (from Lemma 5.2), if one views $\delta$ as a small constant, we recover the same asymptotics with $d+I$ in place of $d$. This is an analog of our main result if one interprets $d+I$ as an “effective VC dimension”.
@@
$\mathrm{Rad}_{\mathcal{D}}(\mathcal{H})=\frac{1}{m}\mathbb{E}_{\boldsymbol{\sigma}\sim\{\pm 1\}^{m}}\left[\sup_{h\in\mathcal{H}}\sum_{i=1}^{m}\sigma_{i}h(x_{i})\right]$ (14)
Theorem C.2.
@@
The above bound again illustrates that incorporating mutual information into the choice of function class yields generalization bounds which decay gracefully with the value of $I$. In the case that each function class has at most $N$ functions, applying Massart’s lemma (Shalev-Shwartz & Ben-David, 2014) yields
@@
C.3 Proof of Theorem C.1 (VC Dimension)
@@
Here the first inequality is Theorem 6.11 from (Shalev-Shwartz & Ben-David, 2014) while the second is Sauer’s lemma. Our strategy will be to argue that the bound above holds with high probability for fixed $\mathcal{H}$ by the bounded differences inequality and conclude using Lemma 5.2.
Proof of Theorem C.1.
@@
i.e. $X_{t}$ is $(1/2m)$-subgaussian. Applying Lemma 5.2 now implies that with probability $1-\delta$, one has
@@
C.4 Proof of Theorem C.2 (Rademacher Complexity)
@@
This is Lemma 26.2 of (Shalev-Shwartz & Ben-David, 2014). Our strategy is equivalent to the above.
@@
We once more define $Y_{t}=\sup_{h\in\mathcal{H}_{t}}\mathrm{gen}_{\mathcal{D}}(h)$ and $X_{t}=Y_{t}-\mathbb{E}[Y_{t}]$. Note that $X_{t}$ changes by at most $2b/m$ when a given datapoint is changed, and thus McDiarmid’s bounded differences inequality again implies that $X_{t}$ is $(2b^{2}/m)$-subgaussian. Then applying Lemma 5.2 yields that with probability $1-\delta$,
@@
Applying Lemma 26.2 (Shalev-Shwartz & Ben-David, 2014) above, we obtain the following almost sure inequality:
@@
We again note that $\mathrm{Rad}_{\mathcal{D}}(\ell\circ\mathcal{H}_{W})$ changes by at most $2b/m$ when changing an individual datapoint. Hence, once centered, it is again $(2b^{2}/m)$-subgaussian. Applying Lemma 5.2 again produces that with probability $1-\delta$,
@@
Noting that $\ell$ is Lipschitz in its first argument, we apply Talagrand’s contraction lemma (see e.g. (Ledoux & Talagrand, 2013; Shalev-Shwartz & Ben-David, 2014)) to conclude the proof of Theorem C.2. ∎